A class consists of 15 boys and 12 girls. They want to elect officers so that the president and secretary are girls, and the vice president and treasurer are boys.
Answer #1

Ans:

As order matter here so,we will use permutations.

Number of ways to elect president and secretary=12P2=12*11=132

Number of ways to elect president and treasurer=15P2=15*14=210

So,Number of ways to elect officers=132*210=27720